OpenStudy (falling_in_katt):

***WILL MEDAL*** divide. \[\frac{ 8y ^{2}-4y+1 }{ 2y-1 }\]

Mehek (mehek14):

divide 8/2 and subtract the exponents \[y ^{2-1}\]

Mehek (mehek14):

@Falling_In_Katt ?

OpenStudy (alekos):

I would have thought that long division was in order

OpenStudy (falling_in_katt):

4y^1?

Mehek (mehek14):

you can just put \(4y\)

Mehek (mehek14):

now -4/2

OpenStudy (falling_in_katt):

-2

Mehek (mehek14):

and \(\dfrac{y}{y}\)

Mehek (mehek14):

\(y^{1-1}\)

OpenStudy (falling_in_katt):

y^0

Mehek (mehek14):

so y cancels out \(4y-2\)

Mehek (mehek14):

last step \(\dfrac{-1}{1}\)

Mehek (mehek14):

actually the other way \(\dfrac{1}{-1}\) doesn't really matter

OpenStudy (falling_in_katt):

-1

Mehek (mehek14):

so your answer is \(4y-2-1\)

OpenStudy (falling_in_katt):

oh okay. thank you!

Mehek (mehek14):

combine -2 - 1 \(4y-3\)

Mehek (mehek14):

and yw

OpenStudy (alekos):

No way. That's not right

OpenStudy (alekos):

The correct answer using polynomial long division is 4y + 1/(2y-1)